accepted. Be sure to show all of your work in a clear and organized manner to earn full credit. Standard Physics/Math Equation Calculations with units Final answer with units 350 grams of solid iron (c 470 J/ kg C) is initially at 200 C and is added to 110 grams of water (c 4187 J/kg C) at 10 C that is inside of a well insulated container. Determine the final temperature of the system upon reaching equilibrium. You may assume that there is not heat loss to the outside air.
Answer #1

Heat lost by iron =Heat gained by Water

M_{Fe}c_{Fe}(200-T)=M_{w}c_{w}(T-10)

0.35*470*(200-T) = 0.11*4187*(T-10)

200-T = 2.8(T-10)

200-T = 2.8T-28

1.8T = 172

T=95.566oC